OpenStudy (anonymous):

Question 1.1. Which decimal is equivalent to ? (Points : 1) 3.917917917… 3.9171717… 3.9177777… 3.91739173917…

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

Equivalent to what?

OpenStudy (anonymous):

help please

OpenStudy (amoodarya):

do like this a= 3.9177777… multiply by 100 100a= 391.77777… now multiply by 10 1000a=3917.7777... 1000a-100a=3917-391 so 900a=3526 a= 3.9177777… =3526/900

OpenStudy (anonymous):

so its a

OpenStudy (amoodarya):

I did not see the choices !

OpenStudy (amoodarya):

b=3.917917917… 1000b=3917.917917917.. now 1000b-b 999b =3917-3 b=3914/999

OpenStudy (anonymous):

Question 1.1. Which decimal is equivalent to ? (Points : 1) http://static.k12.com/eli/ecollege/593/2_31561/2_43701_1_31680/3d4673466fa75aa3c08e24b2abb9a2ee414a2c1e/media/b6d249d4698d8d0ab16c49d38e0f114b2b2a1308/mathml_59c63a0e5f5b83f5a45f886666f74d676ddd3f5c_1.gif A) 3.917917917… B)3.9171717… C)3.9177777… D)3.91739173917…

OpenStudy (anonymous):

help!!!!!!!!!!! please

OpenStudy (amoodarya):

-- 3.9 17=3.9171717171717171717.. --- 3.917=3.917917917917... - 3.917=3.91777777777777777..